Two particles A and B having charges of + 2.00 × 10-6 C and of -4.00 × 10-6 C respectively are held fixed at a separation of 20.0 cm. Locate the point(s) on the line AB where (a) the electric field is zero (b) the electric potential is zero.
44. A circular wire-loop of radius α carries a total charge Q distributed uniformly over its length. A small length dL of the wire is cut off. Find the electric field at the centre due to the remaining wire.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. q base 2 = 2 × 10^–6 C, q base 1^2 = – 4 × 10^–6 C, r = 20 cm = 0.2 m (E base 1 = electric field due to q base 1, E base 2 = electric field due to q base 2) ⇒ (r - x)^2/x^2 = -q base 2/q base 1 ⇒ (r - 1)^2/x = -q^2/q base 1 = 4 * 10^-6/2 * 10^-6 = 1/2 ⇒ (r/x - 1) = 1/√2 = 1/1.414 ⇒ r/x = 1.414 + 1 = 2.414 ⇒ x = r/2.414 = 20/2.414 = 8.285 cm